OpenStudy (anonymous):

Kevin and Conner tracked their exercise last week. Kevin ran fourteen laps around the school track, 8 laps around his neighborhood, and 3 miles on the treadmill. Conner ran 7 laps around the school track and eleven miles on the treadmill. Write and simplify an expression for the total combined distance that Kevin and Conner ran. Show your work. Use t for the distance in laps around their school track and n for the distance in laps around Kevin’s neighborhood.

OpenStudy (bibby):

You can represent their running in equations and add the two but I'm not sure how to proceed. kevin = 14t + 8n + 3 = 25 conner = 7t + 0n + 11 = 18 I guess you can add the two and get 21t + 8n + 14 = 43 21t + 8n = 29 Tag some peoples and keep bumping

OpenStudy (anonymous):

Thanks ^_^

OpenStudy (anonymous):

HOORAH ヽ(*・ω・)ノ